OpenStudy (anonymous):

Evaluate the expression: v ⋅ w Given the vectors: r = <5, -5, -2>; v = <2, -8, -8>; w = <-2, 6, -5> Helppppp :_(

OpenStudy (jdoe0001):

well, the same as you got the dot product of the 2D vectors

OpenStudy (anonymous):

wait what?

OpenStudy (jdoe0001):

<a, b, c> \(\cdot <d, e, f> => a*d + b*e + c*f

OpenStudy (jdoe0001):

<a, b, c> \(\bf\large \cdot \) <d, e, f> => a*d + b*e + c*f

OpenStudy (anonymous):

vector v=2i-8j-8k vector w=-2i+6j-5k v .w =(2i-8j-8k).(-2i+6j-5k) \[=\left( 2 \right)\left( -2 \right)+\left( -8 \right)\left( 6 \right)+\left( -8 \right)\left( -5 \right)\] now you can calculate.

OpenStudy (anonymous):

ahhh okay I see so -12!?

OpenStudy (anonymous):

@surjithayer and @jdoe0001

OpenStudy (anonymous):

you guys still there??

OpenStudy (jdoe0001):

5*2*-2 => -20 -5*-8*6 => +240 -2*-8*-5 => -80 (-20) + (240) + (-80)

OpenStudy (jdoe0001):

at times the system gets a bit lagged :/

OpenStudy (anonymous):

that's alright :) and oh okay so 140? sorry before I did what @surjithayer said so I got -12?

OpenStudy (anonymous):

But it should be 140 right?

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

ohhh wait

OpenStudy (jdoe0001):

I didn't do only u\(\cdot\) w

OpenStudy (jdoe0001):

\(\bf u\cdot w will be 5*-2 = > -10 -5* 6 => -30 -2*-5 => +10 so (-10) + (-30) + (10)

OpenStudy (jdoe0001):

\(\bf u\cdot w\) will be 5*-2 = > -10 -5* 6 => -30 -2*-5 => +10 so (-10) + (-30) + (10)

OpenStudy (anonymous):

so -30?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

wait so why did we have to find 140??

OpenStudy (jdoe0001):

that was a mistake I made, multiplying over too many vecors :/

OpenStudy (anonymous):

ahh okay so v*u = -30 :)

OpenStudy (anonymous):

thank you so much again :)

OpenStudy (jdoe0001):

yw